the sum of the squares of two consecutive numbers is 61.What are those numbers
Found 2 solutions by Boreal, mananth:
Answer by Boreal(15235)

x^2+(x+1)^2=61
x^2+x^2+2x+1=61
so 2x^2+2x-60=0
so x^2+x-30=0
(x+6)(x-5)=0
x=-6; and -6 and -5 have squares who add to 61.
x=5, and 5 and 6 do the same
Those are the two pairs of numbers.

Answer by mananth(16946)

Let the consecutive numbers be x, x+1
x^2 +(x+1)^2 =61
x^2 +x^2 +2x +1 =61
2x^2 +2x -60 =0
divide by 2
x^2 +x -30=0
(x+6)(x-5) =0
x= -6 Or x=5
If one number is -6 other number is -5
If one number is 5 other number is 6